# Structure-Aware Compound-Protein Affinity Prediction via Graph Neural Network with Group Lasso Regularization thanks: Code available: https://github.com/FrancisShizy/SAGGLR

### v1

@@ Abstract
Artificial intelligence accelerates drug discovery via learning drug structure information and molecular representation. Explainable artificial intelligence approaches have been increasingly applied in such areas to rationalize drug property prediction. However, building end-to-end explainable machine learning models for structure-activity relationship (SAR) modeling for compound property prediction faces many challenges, such as the limited number of compound-protein interaction activity data for one specific protein target, and plenty of subtle changes in molecular configuration sites significantly affecting molecular properties. In this context, we take advantage of molecule pairs with activity cliffs that have common scaffolds and distinctive substituent sites with large differences in potency for specific protein targets. We developed a deep learning approach by utilizing the activity cliff-based data of paired molecules targeting three proto-oncogene tyrosine-protein kinase Src proteins (i.e., PDB IDs include 1O42, 2H8H, and 4MXO) respectively. Specifically, we implemented graph neural network (GNN) methods to obtain atom-level feature information and predict compound-protein affinity (i.e., half maximal inhibitory concentration, IC50). In addition, we trained GNN models with different structure-aware loss functions to adequately leverage molecular property and structure information. We also utilized group lasso and sparse group lasso to prune and highlight molecular subgraphs and enhance the structure-specific model explainability for the predicted property difference in molecular activity-cliff pairs. We improved drug property prediction by integrating common and uncommon node information and using sparse group lasso, reducing the average root mean squared error (RMSE) by 12.70%, and achieving the lowest averaged RMSE=0.2551 and the highest PCC=0.9572. Furthermore, applying group lasso and sparse group lasso enhances feature attribution methods that estimate the contribution of each atom in the molecular graphs by boosting global direction scores and atom-level accuracy in atom coloring prediction, which improves model interpretability in drug discovery pipelines, particularly in investigating important molecular substructures in lead optimization.
@@ 1 Introduction
In this paper, we developed a novel framework by leveraging the MMP-cliff in the training to discern key moiety structures related to the compound-protein affinity. We also modified the traditional loss function by adding regularization terms using group lasso and sparse group lasso to improve explainability. We mainly applied our framework on predicting compound affinity for Src [16], an important protein kinase that is important for many diseases such as Alzheimer’s disease [17] and cancers [18]. Our study leads to the findings below: (i) Deep learning model performance for drug property prediction is improved by using graph neural networks by capturing and integrating common and uncommon substructure information of activity-cliffs pairs in the model. The improvement also takes advantage of the loss functions with regularization to prune and highlight molecular subgraphs. (ii) Applying regularization methods such as group lasso and sparse group lasso enhance feature attribution by boosting global direction scores and atom-level accuracy in atom coloring prediction. They provide stable subgraph-level model explainability for molecular property prediction. (iii) Our approach can enhance the interpretability of models in drug discovery and virtual screening pipelines, particularly in investigating important molecular substructures in lead optimization. We introduce the rationales of our framework below in details.
@@
Drug-protein interaction (DPI) modeling tasks, including compound-protein affinity prediction and functional group-based drug property prediction and classification, are important research areas for traditional SAR models and graph-level tasks of deep learning models, such as GNN models. These models have been applied to preclinical drug discovery, virtual screening, and lead optimization [2] [19]. Instead of directly inputting imbalanced drug-protein interaction data into such models, one strategy is to combine molecular activity information such as half-maximal inhibitory concentration (IC50) with structure information (e.g., scaffolds and motifs) to integrate datasets. The loss functions for the models are defined based on molecular similarity and potency difference criteria. Activity cliffs (ACs) are generally defined as pairs or groups of structurally similar compounds that are active against the same target but have large differences in potency [20]. A matched molecular pair (MMP) is a pair of compounds that share a common core structure and are distinguished by substituents at a single site [21]. MMP-based AC (MMP-cliff) can be exploited to capture a large difference in potency between the participating compounds by applying the maximum common substructure (MCS) formalism [22] to calculate the molecular scaffold between pairs of compounds binding to a specific target. For two compounds to be considered as sharing a molecular scaffold, the shared substructure must constitute at least a specified fraction of each molecule’s framework (i.e., percentage thresholds of shared atoms between the matched pairs). Thus, such activity-cliff data can provide critical information regarding key structural components that determine the drastic difference in MMPs. They can then be applied to predict drug-protein binding affinities and provide topology-specific model explainability for drug property differences for specific protein targets. In many previous studies [21][22][23], models were built by training on compound-protein interaction datasets, including various targeting proteins downloaded from cheminformatics databases, to predict compound-protein affinity. While these models achieved reasonable average accuracy over a large number of targets, the accuracy on individual protein targets is often unsatisfactory. By exploiting the information offered by MMP-cliff, we can focus our training on data for a specific protein target.
@@ 2 Methods
Figure 1: Model structure illustration. Considering a pair of compounds $c_{i}$ and $c_{j}$ that share a scaffold in the red circle and decorations in the green circle, we used GNN and MPNN to learn latent node representations for both common and uncommon nodes. Such node-level information was aggregated to predict graph-level drug-protein binding affinity, and then the mean squared errors for predicted and experimental ones were calculated. Additionally, masking functions for common and uncommon nodes and readout functions were applied to normalize the information, and multilayer perceptrons (MLPs) were combined with group lasso and sparse group lasso to create subgraph node loss for both common and uncommon nodes. The two loss functions were minimized during the training process. "CN" for common nodes and structure, "UCN" for uncommon nodes and structure, "N" for both common and uncommon nodes.
@@ 3.1 Model performance
We assessed model performance under different combinations of loss functions and regularization schemes. Table 1 compares model prediction performance by using averaged root mean square error (RMSE) and Pearson correlation coefficient (PCC) values and their weighted values based on the testing set. We evaluated RMSE and PCC on all three test datasets, first by assigning equal weight to each dataset and then by weighting them according to the count of active-cliff pairs in their respective test sets. We defined the MSE loss on the absolute predicted versus experimental binding affinities of molecule pairs, increased loss function complexity from only uncommon node information to both common and uncommon substructures, and added group lasso or sparse group lasso regularization to such loss functions. They have an increasingly positive influence on model performance, reflected by the decreasing trend of averaged RMSE values and the increasing trend of averaged PCC values. The model with the loss function considering both common and uncommon substructures and applying sparse group lasso performs the best compared with other loss function settings, which achieves the lowest averaged $\text{RMSE}=0.2551$ and the highest averaged $\text{PCC}=0.9572$. Moreover, drug property prediction was improved by integrating common and uncommon node information and using regularization, reducing the average root mean squared error (RMSE) by 11.19% (with group lasso) and 12.70% (with sparse group lasso) compared to the loss function only considering the uncommon nodes without penalties.
Table 1: Model Performance of different loss function settings.
Loss avg. RMSE avg. PCC w. RMSE w. PCC
${}^{\mathrm{a}}L_{\text{UCN}}$ 0.5037 0.8425 0.5444 0.8580
${}^{\mathrm{b}}L_{\text{N}}$ 0.2922 0.9425 0.2623 0.9669
${}^{\mathrm{c}}L_{\text{N+GL}}$ 0.2595 0.9542 0.2490 0.9693
${}^{\mathrm{d}}L_{\text{N+SGL}}$ 0.2551 0.9572 0.2422 0.9716
†‡Averaged root mean square error (RMSE) and Pearson correlation
coefficient (PCC) cross three datasets equally and weight-averaged
by their number of active-cliff pairs in testing sets.
aLoss function only for uncommon substructure (same setting as
Kenza Amara et al. [31])
bLoss function for both common and uncommon substructure.
cLoss function with group lasso.
dLoss function with sparse group lasso.
@@
Figure 2: Model performance for the molecular targeting three kinases respectively in testing sets under different loss function settings. It shows the consistent trends of RMSE decreasing and PCC increasing as more node information (from only uncommon nodes to both common and uncommon nodes) and regularization methods were added (from no penalty items for loss functions to adding group lasso and sparse group lasso). This suggests that the models using loss functions with regularization perform better on affinity prediction.
@@ 3.2 Model Evaluation
We evaluated and compared the averaged global direction scores for $L_{MSE}+L_{N}$ with and without group lasso by scatter plots with connecting lines for the four feature attribution approaches respectively, using compounds targeting three kinases in the testing sets. The used compound pairs are considered at the minimum 50% MCS threshold from 50% to 95% in 5% increments. Figure 3 suggests a relationship where the global direction scores predicted by the model with the loss function incorporating group lasso (y-axis) tend to exceed those of the model without the penalty item (x-axis) across the testing dataset under different minimum common substructure thresholds. It shows that there is an increase of 47.13% in the global direction score for the feature attribution CAM with a significant Wilcoxon test p-value of 0.0002 (Figure 3A), an increase of 14.83% in the global direction score for Grad-CAM with a significant Wilcoxon test p-value of 0.0059 (Figure 3B), an increase of 15.4% in the global direction score for Gradient $\times$ Input with a significant Wilcoxon test p-value of 0.002 (Figure 3C), and an increase of 8.49% in the global direction score for IG with a significant Wilcoxon test p-value of 0.0098 (Figure 3D). Supplement figures (Figure 5, 6, and 7) also include scatter plots with connecting lines for compounds targeting three kinases respectively, reflecting the increase of global direction scores for $L_{MSE}+L_{N}$ with group lasso compared to the ones without group lasso. The results suggest that introducing regularization like group lasso improves the performance of feature attribution methods.
Figure 3: Comparison of averaged global direction scores for $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ with and without group lasso by scatter plot with connecting lines. The plot was used for comparing graph-level global direction scores to show the distribution difference of averaged predicted global direction values with $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ (x-axis) and $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ with group lasso (y-axis) loss functions under different minimum common substructure thresholds. Compound pairs are considered at the minimum 50% MCS threshold and from 50% to 100% in 5% increments. The text box reports the increased percentage of global direction scores with group lasso regularization for all the feature attribution including CAM (Figure. 3A) having 47.13% increase and a significant Wilcoxon test p-value of 0.0002; Grad-CAM (Figure. 3B) having 14.83% increase and p-value of 0.0059; Gradient $\times$ Input (Figure. 3C) having 15.4% increase and p-value of 0.002; IG (Figure. 3D) having 8.49% increase and p-value of 0.0098.
@@ Acknowledgment
Supplementary figures for global direction scores

### v2

@@ Abstract
Artificial intelligence accelerates drug discovery via learning drug structure and molecular representation. Explainable artificial intelligence (XAI) approaches are increasingly applied in such areas to rationalize drug property prediction. However, building end-to-end explainable models for structure-activity relationship (SAR) modeling for compound property prediction faces many challenges, such as the limited number of compound-protein interaction activity data for specific protein targets, and plenty of subtle changes in molecular configuration sites significantly affecting molecular properties. Thus, optimally leveraging structural and property information and identifying key moieties related to compound–protein affinity for specific targets is essential. In this context, we exploit pairs of molecules with activity cliffs that share scaffolds but differ at substituent sites, characterized by large potency differences for specific protein targets. We propose a framework by implementing graph neural networks (GNNs) to leverage property and structure information from activity cliff pairs to predict compound-protein affinity (i.e., half maximal inhibitory concentration, IC50). To enhance model performance and explainability, we train GNNs with structure-aware loss functions using group lasso and sparse group lasso regularizations, which prune and highlight molecular subgraphs relevant to activity differences. We applied this framework to activity cliff data of molecules targeting three proto-oncogene tyrosine-protein kinase Src proteins (PDB IDs: 1O42, 2H8H, 4MXO). Our approach improved property prediction by integrating common and uncommon node information with sparse group lasso, as reflected in reduced root mean squared error (RMSE) and improved Pearson’s correlation coefficient (PCC). Applying regularizations also enhances feature attribution for GNN by boosting graph-level global direction scores and improving atom-level coloring accuracy. These advances strengthen model interpretability in drug discovery pipelines, particularly for identifying critical molecular substructures in lead optimization.
@@ 1 Introduction
In this paper, we developed a novel framework by leveraging the matched molecular pair(MMP)-based cliff in the training to discern key moiety structures related to the compound-protein affinity. We also modified the traditional loss function by adding regularization terms using group lasso and sparse group lasso to improve explainability. We mainly applied our framework on predicting compound affinity for Src [16], an important protein kinase that is important for many diseases such as Alzheimer’s disease [17] and cancers [18]. Our study leads to the findings below: (i) Deep learning model performance for drug property prediction is improved by using graph neural networks by capturing and integrating common and uncommon substructure information of activity-cliffs pairs in the model. The improvement also takes advantage of the loss functions with regularization to prune and highlight molecular subgraphs. (ii) Applying regularization methods such as group lasso and sparse group lasso enhance feature attribution by boosting global direction scores and atom-level accuracy in atom coloring prediction. They provide stable subgraph-level model explainability for molecular property prediction. (iii) Our approach can enhance the interpretability of models in drug discovery and virtual screening pipelines, particularly in investigating important molecular substructures in lead optimization. We introduce the rationales of our framework below in details.
@@
Drug-protein interaction (DPI) modeling tasks, including compound-protein affinity prediction and functional group-based drug property prediction and classification, are important research areas for traditional SAR models and graph-level tasks of deep learning models, such as GNN models. These models have been applied to preclinical drug discovery, virtual screening, and lead optimization [2] [19]. Instead of directly inputting imbalanced drug-protein interaction data into such models, one strategy is to combine molecular activity information such as half-maximal inhibitory concentration (IC50) with structure information (e.g., scaffolds and motifs) to integrate datasets. The loss functions for the models are defined based on molecular similarity and potency difference criteria. Activity cliffs (ACs) are generally defined as pairs or groups of structurally similar compounds that are active against the same target but have large differences in potency [20]. A matched molecular pair (MMP) is a pair of compounds that share a common core structure and are distinguished by substituents at a single site [21]. MMP-based AC (MMP-cliff) can be exploited to capture a large difference in potency between the participating compounds by applying the maximum common substructure (MCS) formalism [22] to calculate the molecular scaffold between pairs of compounds binding to a specific target (Figure 1). For two compounds to be considered as sharing a molecular scaffold, the shared substructure must constitute at least a specified fraction of each molecule’s framework (i.e., percentage thresholds of shared atoms between the matched pairs). Thus, such activity-cliff data can provide critical information regarding key structural components that determine the drastic difference in MMPs. They can then be applied to predict drug-protein binding affinities and provide topology-specific model explainability for drug property differences for specific protein targets. In many previous studies [21][22][23], models were built by training on compound-protein interaction datasets, including various targeting proteins downloaded from cheminformatics databases, to predict compound-protein affinity. While these models achieved reasonable average accuracy over a large number of targets, the accuracy on individual protein targets is often unsatisfactory. By exploiting the information offered by MMP-cliff, we can focus our training on data for a specific protein target.
Figure 1: An illustration for an example pair of molecules targeting Src kinase 1O42 with activity cliffs ($\Delta pK_{i}=3.85$). The paired two molecules share common and uncommon substructures. Common substructures consist of uncolored nodes (atoms) and edges (bonds), and uncommon substituents consist of colored nodes and edges (blue & red circles for sites in the pair)
@@ 2 Methods
Figure 2: Model structure illustration. Considering a pair of compounds $c_{i}$ and $c_{j}$ that share a scaffold in the red circle and decorations in the blue circle, GNN and MPNN were applied to learn latent node representations for both common and uncommon nodes. Such node-level information was aggregated to predict graph-level drug-protein binding affinity, and then the mean squared errors for predicted and experimental ones were calculated. Additionally, masking functions for common and uncommon nodes and readout functions were applied to normalize the information, and multilayer perceptrons (MLPs) were combined with group lasso and sparse group lasso to create subgraph node loss for both common and uncommon nodes. The two loss functions were minimized during the training process. "CN" for common nodes and structure, "UCN" for uncommon nodes and structure, "N" for both common and uncommon nodes. Regularization methods were also applied to prune and select node-level information in activity cliff pairs. Group lasso considers only group-level sparsity (nodes within the same subgroup having the same color depth representing the same weights after pruning), and sparse group lasso considers both group-level and within-group sparsity (nodes within the same subgroup having different color depth representing various weights after pruning).
@@ 3.1 Model performance
We assessed model performance under different combinations of loss functions and regularization schemes via 5-fold cross validation. Table 1 compares model prediction performance by using averaged root mean square error (RMSE) and Pearson correlation coefficient (PCC) values and their weighted values based on the testing set. We evaluated RMSE and PCC on all three test datasets, first by assigning equal weight to each dataset and then by weighting them according to the count of active-cliff pairs in their respective test sets. We defined the MSE loss on the absolute predicted versus experimental binding affinities of molecule pairs, increased loss function complexity from only uncommon node information to both common and uncommon substructures, and added group lasso or sparse group lasso regularization to such loss functions. They have an increasingly positive influence on model performance, reflected by the decreasing trend of averaged RMSE values and the increasing trend of averaged PCC values. The model with the loss function considering both common and uncommon substructures and applying sparse group lasso performs the best compared with other loss function settings, which achieves the lowest averaged $\text{RMSE}=0.2143$ and lowest weighted averaged $\text{RMSE}=0.1984$, and the highest averaged $\text{PCC}=0.9615$ and lowest weighted averaged $\text{PCC}=0.9468$. Moreover, drug property prediction was improved by integrating common and uncommon node information and using regularization, reducing the averaged RMSE by 15.71% and the weighted averaged RMSE by 11.62% (with group lasso), and decreasing the averaged RMSE by 17.48% and the weighted averaged RMSE by 11.62% (with sparse group lasso) compared to the loss function only considering the uncommon nodes without penalties.
Table 1: Model performance of different loss functions (5-fold cross-validation)
[…]
$L_{\text{UCN}}$a 0.2597 ($\pm$0.0254) 0.9541 ($\pm$0.0096) 0.2245 ($\pm$0.0303) 0.9384 ($\pm$0.0164)
$L_{\text{N}}$b 0.2371 ($\pm$0.0221) 0.9564 ($\pm$0.0056) 0.2143 ($\pm$0.0278) 0.9386 ($\pm$0.0161)
$L_{\text{N+GL}}$c 0.2189 ($\pm$0.0219) 0.9609 ($\pm$0.0071) 0.1984 ($\pm$0.0205) 0.9468 ($\pm$0.0146)
$L_{\text{N+SGL}}$d 0.2143 ($\pm$0.0185) 0.9615 ($\pm$0.0078) 0.1984 ($\pm$0.0150) 0.9466 ($\pm$0.0140)
  • $\mathrm{\dagger\ddagger}$
    Averaged root mean square error (avg. RMSE) and Pearson correlation coefficient (avg. PCC) across three datasets equally, and weighted average using their number of active-cliff pairs in testing sets (w. RMSE & w. PCC)after 5-fold cross-validation.
  • a
    Loss function only for uncommon substructure (same setting as Kenza Amara et al. [31]).
  • b
    Loss function for both common and uncommon substructure.
  • c
    Loss function with group lasso.
  • d
    Loss function with sparse group lasso.
@@
Figure 3: Model performance for the molecules in testing sets targeting three kinases respectively, under different loss function settings via 5-fold cross-validation. Numbers for each bar plot indicate the averaged RMSE or PCC values for each loss function setting, and error bars represent standard deviation. It shows the consistent trends of RMSE decreasing and PCC increasing as more node information (from only uncommon nodes to both common and uncommon nodes) and regularization methods were added (from no penalty items for loss functions to adding group lasso and sparse group lasso). This suggests that the models using loss functions with regularization perform better on affinity prediction.
@@ 3.2 Model Evaluation
We evaluated and compared the averaged global direction scores for $L_{MSE}+L_{N}$ with and without group lasso by scatter plots with connecting lines for the four feature attribution approaches respectively, using compounds targeting three kinases in the testing sets. The used compound pairs are considered at the minimum 50% MCS threshold from 50% to 95% in 5% increments. Figure 4 suggests a relationship where the global direction scores predicted by the model with the loss function incorporating group lasso (y-axis) tend to exceed those of the model without the penalty item (x-axis) across the testing dataset under different minimum common substructure thresholds. It shows that there is an increase of 47.13% in the global direction score for the feature attribution CAM with a significant Wilcoxon test p-value of 0.0002 (Figure 4A), an increase of 14.83% in the global direction score for Grad-CAM with a significant Wilcoxon test p-value of 0.0059 (Figure 4B), an increase of 15.4% in the global direction score for Gradient $\times$ Input with a significant Wilcoxon test p-value of 0.002 (Figure 4C), and an increase of 8.49% in the global direction score for IG with a significant Wilcoxon test p-value of 0.0098 (Figure 4D). Supplement figures (Figure 6, 7, and 8) also include scatter plots with connecting lines for compounds targeting three kinases respectively, reflecting the increase of global direction scores for $L_{MSE}+L_{N}$ with group lasso compared to the ones without group lasso. The results suggest that introducing regularization like group lasso improves the performance of feature attribution methods.
Figure 4: Comparison of averaged global direction scores for $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ with and without group lasso by scatter plot with connecting lines. The plot was used for comparing graph-level global direction scores to show the distribution difference of averaged predicted global direction values with $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ (x-axis) and $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ with group lasso (y‑axis) loss functions under different minimum common substructure thresholds. Compound pairs are considered at the minimum 50% MCS threshold and from 50% to 100% in 5% increments. The text box reports the increased percentage of global direction scores with group lasso regularization for all the feature attribution including CAM (Figure. 4A) having 47.13% increase and a significant Wilcoxon test p-value of 0.0002; Grad-CAM (Figure. 4B) having 14.83% increase and p-value of 0.0059; Gradient $\times$ Input (Figure. 4C) having 15.4% increase and p-value of 0.002; IG (Figure. 4D) having 8.49% increase and p-value of 0.0098.
@@ Acknowledgment
Supplementary